17 (a) Prove that a permutation π in the Permutation Cipher is an involutory key if and only if π(i) = j implies π(j) = i, for all i, j E {1, . . . , m}

(b) Determine the number of involutory keys in the Permutation Cipher for m = 2,3, 4, 5 and 6.

0 0

> For a given permutation cipher “DEEEGLLMMOOOVVY”, please decrypt it (and show the details). how to solve it ?

Naqed Shaker Tue, Nov 23, 2021 1:41 AM

Add a comment Improve this question Transcribed image text
Answer #1

Solution 0 since a key k is invoulogy is Eu (0) : Dx (d) foo any plainted a. In the case of peomutation cipheo, this oelalive0 let Pm denote the number of involubony brey of block : size m. we have P=1 and P2 =2 Now, suppose had we have computed thePm Pm-1 + (m-1) Pm-2 Foo m upto 6 we have that P,= 1; P2=2 Then Pz= Patapi P₂=2+2(1) Pu= P₂ +382 : 4+3127 = 4+6=10 Pg= putuP3Dear student if you have any Doubt Please Ask, Thank you.